### v2

# A generalization of the Romanoff theorem

Abstract.
Let $\mathcal{P}$ be the set of primes and $\mathbb{N}$ the set of positive integers. Let also $r_{1},...,r_{t}$ be positive real numbers and $R_{2}(r_{1},...,r_{t})$ the set of odd integers which can be represented as
$p+2^{\lfloor k_{1}^{r_{1}}\rfloor}+\cdot\cdot\cdot+2^{\lfloor k_{t}^{r_{t}}\rfloor},$
where $p\in\mathcal{P}$ and $k_{1},...,k_{t}\in\mathbb{N}$. Recently, Chen and Xu proved that the set $R_{2}(r_{1},...,r_{t})$ has positive lower asymptotic density, provided that $r_{1}^{-1}+\cdot\cdot\cdot+r_{t}^{-1}\geq 1$ and at least one of $r_{1},...,r_{t}$ is an integer. Their result reduces to the famous theorem of Romanoff by taking $t=r_{1}=1.$ In this note, we remove the unnecessary condition: ‘at least one of $r_{1},...,r_{t}$ is an integer’.
Key words and phrases:
Romanoff type problems, primes, powers of $2$
2010 Mathematics Subject Classification:
11P32, 11A41, 11B13
ORCID: $0$-0001-7016-309X
1. Introoduction
The aim of this note is the following generalization of the classical theorem named after Romanoff [21], which states that the set of odd integers whose elements can be written as the sum of a prime and a power of two has positive lower asymptotic density.
Theorem 1.
Let $r_{1},...,r_{t}$ be positive real numbers. Then the set
$R_{2}(r_{1},...,r_{t})=\left\{\text{odd}~{}n:n=p+2^{\lfloor k_{1}^{r_{1}}\rfloor}+\cdot\cdot\cdot+2^{\lfloor k_{t}^{r_{t}}\rfloor},~{}p\in\mathcal{P},k_{1},...,k_{t}\in\mathbb{N}\right\}$
has positive lower asymptotic density if and only if $r_{1}^{-1}+\cdot\cdot\cdot+r_{t}^{-1}\geq 1$, where $\mathcal{P}$ and $\mathbb{N}$ are the sets of primes and positive integers, respectively.
This generalization was introduced in a recent article of Chen and Xu [3] but only with an additional condition (which is crucial in their argument): ‘at least one of $r_{1},...,r_{t}$ is an integer’.
Before the proof of Theorem 1, We mention the history involving the Romanoff theorem below. de Polignac [19] conjectured that every odd number greater than 1 is the sum of a prime and a power of 2. But soon, he [19] recognized that Euler had already mentioned a counterexample 959 in a letter to Goldbach. Towards the positive direction, Romanoff [21] proved that there is a positive proportion of the odd numbers which can be written as the sum of a prime and a power of two. Later, Erdős [10] and van der Corput [4] showed that the odd numbers which can not be represented by the form $p+2^{m}$ with $p\in\mathcal{P}$ and $m\in\mathbb{N}$ also possess positive lower asymptotic density. Crocker [5] proved that there are infinitely many odd numbers which cannot be represented as $p+2^{m}+2^{n}$ with $p\in\mathcal{P}$ and $m,n\in\mathbb{N}$, which was later improved significantly by Chen–Feng–Templier [2] and Pan [17].
Erdős paid much attention Romanoff’s theorem and the proof of it was later simplified by Erdős and Turán [11] and Erdős [9]. In [10], Erdős extended powers of $2$ in the Romanoff theorem to a more general sequence $\{a_{k}\}$. Precisely, let $a_{1}<a_{2}<\cdots$ be an infinite sequence of positive integers satisfying $a_{k}|a_{k+1}$. Then Erdős proved that the necessary and sufficient condition that the sequence $p+a_{k}$ should have positive density is that
$\limsup_{k\rightarrow\infty}\frac{\log a_{k}}{k}<\infty\quad\text{and}\quad\sum_{d|a_{i}}\frac{1}{d}<c$
for some absolute constant $c$ and any $a_{i}$.
Following a question of Chen [1], Yang and Chen [25] considered variants of the Romanoff theorem with $c$-condition. A set $\mathcal{B}$ is said to be of $c$-condition if there is some absolute constant $10<c<$ 1 so that $\mathcal{B}(cx)\gg\mathcal{B}(x)$. Yang and Chen studied the following sumset
$\mathcal{S}=\big{\{}p+b:p\in\mathcal{P},b\in\mathcal{B}\big{\}}.$
They proved that
$\frac{x}{\log x}\min\bigg{\{}\mathcal{B}(x),\frac{\log x}{\log\log x}\bigg{\}}\ll\mathcal{S}(x)\ll\frac{x}{\log x}\min\big{\{}\mathcal{B}(x),\log x\big{\}}.$
As an application, Yang and Chen obtained the following estimates
#$\#\big{\{}n\leq x:n=p+2^{a^{2}}+2^{b^{2}},~{}p\in\mathcal{P},a,b\in\mathbb{N}\big{\}}\gg\frac{x}{\log\log x}.$
Moreover, for any $m$ they constructed a subset $\mathcal{B}$ with $c$-condition so that
$\mathcal{B}(x)=\frac{1+o(1)}{m+1}\left(\frac{\log x}{\log\log x}\right)^{m+1}$
and $\mathcal{S}(x)\ll x/\log\log x$. These naturally led them to the following two problems.
Problem 1.
Does there exist a real number $\alpha>0$ and a subset $\mathcal{B}$ of $\mathbb{N}$ with $c$-condition such that $\mathcal{B}(x)\gg x^{\alpha}$ and $\mathcal{S}(x)\ll x/\log\log x$?
Problem 2.
Does there exist a positive integer $k$ such that the set of positive integers which can be represented as $p+\sum_{i=1}^{k}2^{m_{i}^{2}}$ with $p\in\mathcal{P}$ and $m_{i}\in\mathbb{N}$ has a positive lower density? If such $k$ exists, what is the minimal value of such $k$?
In a prior article, the author [6] completely solved Problem 2 by showing that $k=2$ is admissible. Thus,
#$\#\big{\{}n\leq x:n=p+2^{a^{2}}+2^{b^{2}},~{}p\in\mathcal{P},a,b\in\mathbb{N}\big{\}}\gg x.$
In a subsequent article, the author [7] answered Problem 1 negatively and proved that for any subset $\mathcal{B}$ of $c$-condition with $\mathcal{B}(x)\gg x^{\alpha}$ we have
$\mathcal{S}(x)\gg x/\log\log\log x.$
I also pointed out that the above lower bound is sharp via concrete examples. Following the line of Problem 2, Chen and Xu [3] obtained Theorem 1 with additional condition: ‘at least one of $r_{1},...,r_{t}$ is an integer’. And in [3, Remark 1], they commented that ‘We believe that it is unnecessary that at least one of $r_{1},...,r_{t}$ is an integer in our theorem’. It is just our purpose to remove this additional condition which forms a clean generalization (Theorem 1) of the Romanoff theorem.
For more variants on the Romanoff theorem, see e.g. Filaseta–Finch–Kozek [14], Pan–Li [18], Lee [15], Shparlinski–Weingartner [22], Elsholtz–Luca–Planitzer [8].
2. proofs
The main difference of the proof is that when we encounter Eq. (9), we can not remove the floor function directly since $r_{1}$ will not be an integer any more (comparing this with the corresponding argument of Chen and Xu [3, Theorem 3.1]). By an elementary handle we transform Eq. (9) to Eq. (10) and then a result of Erdős and Turán could be used therein. This finite form result on uniform distribution is crucial here since the modulo $e_{2}(d)$ could be significantly greater than the range of $k_{1}$ (see Eq. (9)), which is different to the usual researches on Piatetski-Shapiro like sequences.
Throughout the proof, $\mathcal{S}(x)$ denotes the number of elements of $\mathcal{S}$ not exceeding $x$.
Proof of Theorem 1.
One can also easily note that $R_{2}(r_{1},...,r_{t})$ is of density zero if $r_{1}^{-1}+\cdot\cdot\cdot+r_{t}^{-1}<1$, making the necessary condition valid. So, we only need to prove the sufficient point.
Suppose first that there exists some $1\leq s\leq t$ such that $r_{s}^{-1}\geq 1$, then
$\left\{2^{\lfloor k_{s}^{r_{s}}\rfloor}:k_{s}\in\mathbb{N}\right\}=2^{\mathbb{N}}$
and hence
$\mathcal{P}+2^{\mathbb{N}}\subset R_{2}(r_{1},...,r_{t}).$
In this case, there is nothing to prove for our theorem as it is just the Romanoff theorem.
From now on, we assume that $r_{i}^{-1}<1$ for any $1\leq i\leq t$. So, we would clearly have $t\geq 2$. On observing Chen–Xu’s result, we can also assume without loss of generality that none of $r_{i}$ belongs to integers. The proof begins with the following observation of Chen and Xu [3, Eq. (4.1)]: There exists an integer $s\geq 2$ such that
$\frac{1}{r_{1}}+\cdot\cdot\cdot+\frac{1}{r_{s-1}}<1\leq\frac{1}{r_{1}}+\cdot\cdot\cdot+\frac{1}{r_{s-1}}+\frac{1}{r_{s}},$
from which it follows that there is real number $\lambda\geq 1$ satisfying that
$\frac{1}{r_{1}}+\cdot\cdot\cdot+\frac{1}{r_{s-1}}+\frac{1}{\lambda r_{s}}=1.$ (1)
Let
$\mathcal{A}=\left\{2^{\lfloor k_{1}^{r_{1}}\rfloor}+\cdot\cdot\cdot+2^{\lfloor k_{s-1}^{r_{s-1}}\rfloor}+2^{\left\lfloor\lfloor k_{s}^{\lambda}\rfloor^{r_{s}}\right\rfloor}\right\}.$
It suffices to prove that the set $\mathcal{P}+\mathcal{A}$ has positive lower asymptotic density since
$\left\{2^{\left\lfloor\lfloor k_{s}^{\lambda}\rfloor^{r_{s}}\right\rfloor}:k_{s}\in\mathbb{N}\right\}\subseteq\left\{2^{\lfloor k_{s}^{r_{s}}\rfloor}:k_{s}\in\mathbb{N}\right\}.$
Denote by $r(n)$ the number of representations of $n=p+a$ with $p\in\mathcal{P}$ and $a\in\mathcal{A}$. We note from Chen and Xu [3, Pages 77 and 78] that
$\mathcal{A}(x)\asymp\log x.$ (2)
It is plain that
$\sum_{n\leq x}r(n)\geq\mathcal{A}(x/2)\mathcal{P}(x/2)\gg\log x(x/\log x)\gg x.$
By the Cauchy–Schwarz inequality we have
$\bigg{(}\sum_{n\leq x}r(n)\bigg{)}^{2}\leq\Bigg{(}\sum_{\begin{subarray}{c}n\leq x\\ r(n)\geq 1\end{subarray}}1\Bigg{)}\bigg{(}\sum_{n\leq x}r(n)^{2}\bigg{)}.$
We are going to show that
$\sum_{n\leq x}r(n)^{2}\ll x,$
from which we would get
$\left(\mathcal{P}+\mathcal{A}\right)(x)=\sum_{\begin{subarray}{c}n\leq x,~{}r(n)\geq 1\end{subarray}}1\gg x.$
Expanding and then exchanging the sum, we obtain
$\sum_{n\leq x}r(n)^{2}=\sum_{\begin{subarray}{c}p_{1}+a_{1}=p_{2}+a_{2}\leq x\\ p_{1},p_{2}\in\mathcal{P},~{}a_{1},a_{2}\in\mathcal{A}\end{subarray}}1\leq\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}\leq a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}p_{2}\leq p_{1}\leq x\\ p_{1}-p_{2}=a_{2}-a_{1}\end{subarray}}1.$
For the innermost sum in the right–hand side above, we separate it into two cases according to $a_{1}=a_{2}$ or not. The contribution of the sum for $a_{1}=a_{2}$ can be controlled by
$A(x)\mathcal{P}(x)\ll x$
via Eq. (2) and the prime number theorem. Thus, to complete our proof, it would suffice to prove that
$S(x):=\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}p_{2}<p_{1}\leq x\\ p_{1}-p_{2}=a_{2}-a_{1}\end{subarray}}1\ll x.$ (3)
For $h\neq 0$, let $\pi_{2}(x,h)$ be the number of prime pairs $p$ and $q$ with $q-p=h$ below $x$. It is well known (see e.g. [16, Theorem 7.3]) that
$\pi_{2}(x,h)\ll\frac{x}{(\log x)^{2}}\prod_{p|h}\left(1+\frac{1}{p}\right).$ (4)
Inserting Eq. (4) into Eq. (3) we obtain
$S(x)\ll\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\frac{x}{(\log x)^{2}}\prod_{p|a_{2}-a_{1}}\left(1+\frac{1}{p}\right).$
We are leading to prove
$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\prod_{\begin{subarray}{c}p|a_{2}-a_{1}\\ p>2\end{subarray}}\left(1+\frac{1}{p}\right)\ll(\log x)^{2}.$
Following an idea of Elsholtz et al. [8], we note that
$\prod_{\begin{subarray}{c}p|a_{2}-a_{1}\\ p\geq\log x\end{subarray}}\left(1+\frac{1}{p}\right)<\left(1+\frac{1}{\log x}\right)^{\log x/\log 2}\ll 1$
since the number of prime factors of $|a_{1}-a_{2}|\leq x$ is trivially not exceeding $\log x/\log 2$. Thus, it reduces to show that
$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\prod_{\begin{subarray}{c}p|a_{2}-a_{1}\\ 2<p<\log x\end{subarray}}\left(1+\frac{1}{p}\right)\ll(\log x)^{2},$
which is equivalent to
$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}d|a_{2}-a_{1}\\ 2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\ll(\log x)^{2},$ (5)
where $P^{+}(d)$ denotes the largest prime factor of $d$ with convention $P^{+}(1)=1$ and $\mu(d)$ the usual Möbius function. On noting the definition of $\mathcal{A}$, it is clear that
$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}d|a_{2}-a_{1}\\ 2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\!\!\frac{\mu^{2}(d)}{d}$ $\leq\!\!\sum_{\begin{subarray}{c}k_{i},\widetilde{k_{i}}\leq\left(\frac{\log 2x}{\log 2}\right)^{\frac{1}{r_{i}}}\\ 1\leq i\leq s-1\end{subarray}}\sum_{k_{s},\widetilde{k_{s}}\leq 2\left(\frac{\log 2x}{\log 2}\right)^{\frac{1}{kr_{s}}}}\sum_{\begin{subarray}{c}d|h_{1}+\cdot\cdot\cdot+h_{s-1}+h_{s}-\left(\widetilde{h_{1}}+\cdot\cdot\cdot+\widetilde{h_{s-1}}+\widetilde{h_{s}}\right)\\ \widetilde{h_{1}}+\cdot\cdot\cdot+\widetilde{h_{s-1}}+\widetilde{h_{s}}<h_{1}+\cdot\cdot\cdot+h_{s-1}+h_{s}\leq x\\ 2\nmid d,~{}P^{+}(d)<\log x\end{subarray}}\!\!\frac{\mu^{2}(d)}{d}$
$\ll(\log x)^{2-\frac{1}{r_{1}}}~{}\max_{|g|\leq x}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ 2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv g\!\ ({\rm{mod}}\ d)\end{subarray}}1,$ (6)
where the last estimate comes from Eq. (1) and
$h_{i}=2^{\left\lfloor k_{i}^{r_{i}}\right\rfloor},~{}~{}~{}\widetilde{h_{i}}=2^{\left\lfloor k_{i}^{r_{i}}\right\rfloor}\quad(1\leq i\leq s-1),\quad\quad h_{s}=2^{\left\lfloor\lfloor k_{s}^{\lambda}\rfloor^{r_{s}}\right\rfloor},~{}~{}~{}\widetilde{h_{s}}=2^{\left\lfloor\left\lfloor\widetilde{k_{s}}^{\lambda}\right\rfloor^{r_{s}}\right\rfloor}.$
For a given $g\leq x$, suppose first that there is at most one solution $k_{1}$ to the congruence equation
$2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv g\ ({\rm{mod}}\ d),$ (7)
then for those $d$ the double sums in right–hand side of Eq. (2) are bounded by
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\leq\prod_{2<p<\log x}\left(1+\frac{1}{p}\right)\ll\log\log x$ (8)
which is admissible for our purpose, where the last estimate follows from the famous theorem of Mertens. We are left over to consider the remaining $d$ for which there are at least two solutions $k_{1}$ satisfying Eq. (7). In this case, let $k_{1,0}$ be such a solution. Then we have
$2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv 2^{\left\lfloor k_{1,0}^{r_{1}}\right\rfloor}\ ({\rm{mod}}\ d).$
So we would get
$\left\lfloor k_{1}^{r_{1}}\right\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))$ (9)
for some $10\leq\ell\leq e_{2}(d)-$ 1, where $e_{2}(d)$ is the least positive integer $m$ so that $d|2^{m}-1$. It follows that
$e_{2}(d)f+\ell\leq k_{1}^{r_{1}}<e_{2}(d)f+\ell+1$
for some $f\in\mathbb{Z}$ from Eq. (9) and vice versa, which clearly means that
$\frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}<\frac{\ell+1}{e_{2}(d)},$
where $\{y\}$ stands for the fractional part of $y$. It would be more pleasant to relax the above inequality to
$\frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{\ell+1}{e_{2}(d)}$ (10)
and we shall consider Eq. (10) from now on. We are in a position to employ a ‘finite’ form regarding the classical theorem of Weyl [24] on uniform distribution which was later established by Erdős and Turán [12, Theorem III].
Lemma 1. If $\varphi_{1},\varphi_{2},\cdot\cdot\cdot,\varphi_{n}$ are real satisfying that
$\bigg{|}\sum_{v=1}^{n}e^{ki\varphi_{v}}\bigg{|}\leq\psi(k)\quad(k=1,2,...,m,...),$
then for $m\geq 1$ and arbitrary $0\leq\alpha<\beta\leq 2\pi$ we have
[…]
with a numerical constant $C$.
Now we take $\varphi_{v}=2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\},n=\left\lfloor\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\right\rfloor,\alpha=\frac{2\pi\ell}{e_{2}(d)}$ and $\beta=\frac{2\pi(\ell+1)}{e_{2}(d)}$ in Lemma 1, then there is some absolute constant $C$ so that
$\Bigg{|}\sum_{\begin{subarray}{c}k_{1}\leq\left(\log 2x/\log 2\right)^{1/r_{1}}\\ \frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{(\ell+1)}{e_{2}(d)}\end{subarray}}\!\!\!\!1-\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{e_{2}(d)}\Bigg{|}<C\left(\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{m+1}+\sum_{k=1}^{m}\frac{\psi(k)}{k}\right),$ (11)
where $m\geq 1$ is a parameter to be decided later and $\psi(k)$ is the quantity satisfying
$\bigg{|}\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\}}\bigg{|}\leq\psi(k).$ (12)
To give an appropriate estimate of $\psi(k)$, we use the van der Corput type bounds [23, Theorems 5.9, 5.11 and 5.13].
Lemma 2. Let $f(x)$ be real and have continuous derivates up to the $j$–th order, where $j\geq 2$. Let $\lambda_{j}\leq f^{(j)}(x)\leq h\lambda_{j}$ (or the same for $-f^{(j)}(x)$). Let $b-a\geq 1,J=2^{j-1}$. Then
$\sum_{a<n\leq b}e^{2\pi if(n)}=O\left(h^{2/J}(b-a)\lambda_{j}^{1/(2J-2)}\right)+O\left((b-a)^{1-2/J}\lambda_{j}^{-1/(2J-2)}\right),$
where the constants implied by big-$O$ are independent of $j$.
Let $f(x)=\frac{kx^{r_{1}}}{e_{2}(d)}$ and $r$ be the least integer greater than $r_{1}$. Then
$f^{(r)}(x)=\frac{k}{e_{2}(d)}r_{1}(r_{1}-1)\cdot\cdot\cdot(r_{1}-r+1)x^{r_{1}-r}.$
Considering with $a\leq x<b\leq 2a$, we clearly have
$\frac{kC_{1}}{e_{2}(d)(2a)^{r-r_{1}}}\leq f^{(r)}(x)\leq 2^{r-r_{1}}\frac{kC_{1}}{e_{2}(d)(2a)^{r-r_{1}}},$
where $C_{1}=r_{1}(r_{1}-1)\cdot\cdot\cdot(r_{1}-r+1)$. Employing Lemma 2 we obtain that
$\sum_{a<n\leq b}e^{2\pi i\frac{kn^{r_{1}}}{e_{2}(d)}}$ $\ll_{r_{1}}a\left(\frac{k}{e_{2}(d)a^{r-r_{1}}}\right)^{1/(2R-2)}+a^{1-2/R}\left(\frac{k}{e_{2}(d)a^{r-r_{1}}}\right)^{-1/(2R-2)}$
$\ll a^{1-\frac{r-r_{1}}{2R-2}}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+a^{1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}},$
where $R=2^{r-1}$. The dichotomy of $a$ between the interval $\left[1,\left(\log 2x/\log 2\right)^{1/r_{1}}\right]$ yields
$\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\}}$ $=\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\frac{v^{r_{1}}}{e_{2}(d)}}$
$\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}}.$
Thus, by Eq. (12) we can choose $\psi(k)$ to satisfy that
$\psi(k)\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}},$
from which it follows directly via partial summations that
$\sum_{k=1}^{m}\frac{\psi(k)}{k}\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{m^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}e_{2}(d)^{\frac{1}{2R-2}}.$ (13)
We take $m=\lfloor\log\log x\rfloor$ and then insert Eq. (13) into Eq. (11). Then
$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{(\ell+1)}{e_{2}(d)}\end{subarray}}\!\!\!\!1=\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{e_{2}(d)}+O_{r_{1}}\left(\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{\log\log x}+\sum_{k=1}^{\lfloor\log\log x\rfloor}\frac{\psi(k)}{k}\right).$ (14)
Collecting Eqs. (2), (8), (13) and (14), we conclude from the analysis above that
$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}d|a_{2}-a_{1}\\ 2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\!\!\frac{\mu^{2}(d)}{d}\ll_{r_{1}}(\log x)^{2-\frac{1}{r_{1}}}\log\log x+S_{1}(x)+S_{2}(x)+S_{3}(x)+S_{4}(x)$ (15)
where
$S_{1}(x):=(\log x)^{2}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{de_{2}(d)},\qquad S_{2}(x):=\frac{(\log x)^{2}}{\log\log x}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d},$
$S_{3}(x):=(\log x)^{2-\frac{r-r_{1}}{r_{1}(2R-2)}}(\log\log x)^{{}^{\frac{1}{2R-2}}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{de_{2}(d)^{\frac{1}{2R-2}}}$
and
$S_{4}(x):=(\log x)^{2-\left(\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}\right)}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)e_{2}(d)^{\frac{1}{2R-2}}}{d}.$
By a result of Erdős and Turán [13], we have
$\sum_{2\nmid d}\frac{1}{d(e_{2}(d))^{\varepsilon}}<\infty$ (16)
for any $\varepsilon>0$. Thus, we deduce that $S_{1}(x)\ll(\log x)^{2}$ and
$S_{3}(x)\ll(\log x)^{2-\frac{r-r_{1}}{r_{1}(2R-2)}}(\log\log x)^{{}^{\frac{1}{2R-2}}}\ll(\log x)^{2}.$
By Eq. (8), we also have $S_{2}(x)\ll(\log x)^{2}$. For $2\nmid d$, we have
$2^{\varphi(d)}\equiv 1\ ({\rm{mod}}\ d)$
by the Euler theorem, where $\varphi(\cdot)$ is the Euler totient function. So $e_{2}(d)\leq\varphi(d)\leq d$, from which it follows that
$S_{4}(x)\leq(\log x)^{2-\left(\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}\right)}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d^{1-\frac{1}{2R-2}}}.$
Changing back the sum to product, we get
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d^{1-\frac{1}{2R-2}}}\leq\prod_{2<p<\log x}\left(1+\frac{1}{p^{1-\frac{1}{2R-2}}}\right)=\exp\left(\sum_{2<p<\log x}\log\left(1+\frac{1}{p^{1-\frac{1}{2R-2}}}\right)\right).$
Employing the elementary inequality $\log(1+z)<z$ for any $z>0$, the right–hand side above is
$<\exp\left(\sum_{2<p<\log x}\frac{1}{p^{1-\frac{1}{2R-2}}}\right).$
Using the prime number theorem as well as the partial summations, we have
$\sum_{2<p<\log x}\frac{1}{p^{1-\frac{1}{2R-2}}}\ll\frac{(\log x)^{\frac{1}{2R-2}}}{\log\log x},$
which implies $S_{4}(x)\ll(\log x)^{2}$ on noting that
$\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}>\frac{1}{2R-2}.$
Therefore, the right–hand side of Eq. (15) is bounded by $(\log x)^{2}$ which proves Eq. (5) and hence Eq. (3). ∎
Acknowledgments
The author is supported by National Natural Science Foundation of China (Grant No. 12201544), Natural Science Foundation of Jiangsu Province, China (Grant No. BK20210784), China Postdoctoral Science Foundation (Grant No. 2022M710121).

### v3

@@
 11P32, 11A41, 11B13
-ORCID: $0$-0001-7016-309X
 1. Introoduction
@@
 Crocker [5] proved that there are infinitely many odd numbers which cannot be represented as $p+2^{m}+2^{n}$ with $p\in\mathcal{P}$ and $m,n\in\mathbb{N}$, which was later improved significantly by Chen–Feng–Templier [2] and Pan [17].
-Erdős paid much attention Romanoff’s theorem and the proof of it was later simplified by Erdős and Turán [11] and Erdős [9].
+Erdős paid much attention to Romanoff’s theorem and the proof of it was later simplified by Erdős and Turán [11] and Erdős [9].
 In [10], Erdős extended powers of $2$ in the Romanoff theorem to a more general sequence $\{a_{k}\}$.
@@
 $\frac{x}{\log x}\min\bigg{\{}\mathcal{B}(x),\frac{\log x}{\log\log x}\bigg{\}}\ll\mathcal{S}(x)\ll\frac{x}{\log x}\min\big{\{}\mathcal{B}(x),\log x\big{\}}.$
-As an application, Yang and Chen obtained the following estimates
+As an application, Yang and Chen obtained the following estimate
 #$\#\big{\{}n\leq x:n=p+2^{a^{2}}+2^{b^{2}},~{}p\in\mathcal{P},a,b\in\mathbb{N}\big{\}}\gg\frac{x}{\log\log x}.$
@@
 If such $k$ exists, what is the minimal value of such $k$?
-In a prior article, the author [6] completely solved Problem 2 by showing that $k=2$ is admissible. Thus,
+In a prior article, the first named author [6] completely solved Problem 2 by showing that $k=2$ is admissible. Thus,
 #$\#\big{\{}n\leq x:n=p+2^{a^{2}}+2^{b^{2}},~{}p\in\mathcal{P},a,b\in\mathbb{N}\big{\}}\gg x.$
-In a subsequent article, the author [7] answered Problem 1 negatively and proved that for any subset $\mathcal{B}$ of $c$-condition with $\mathcal{B}(x)\gg x^{\alpha}$ we have
+In a subsequent article, the first named author [7] answered Problem 1 negatively and proved that for any subset $\mathcal{B}$ of $c$-condition with $\mathcal{B}(x)\gg x^{\alpha}$ we have
 $\mathcal{S}(x)\gg x/\log\log\log x.$
-I also pointed out that the above lower bound is sharp via concrete examples.
+The first named author also pointed out that the above lower bound is sharp via concrete examples.
 Following the line of Problem 2, Chen and Xu [3] obtained Theorem 1 with additional condition: ‘at least one of $r_{1},...,r_{t}$ is an integer’.
@@
 2. Proofs
-The main difference of the proof is that when we encounter Eq.
-(9), we can not remove the floor function directly since $r_{1}$ will not be an integer any more (comparing this with the corresponding argument of Chen and Xu [3, Theorem 3.1]).
-By an elementary handle we transform Eq.
-(9) to Eq.
-(10) and then a result of Erdős and Turán could be used therein.
-This finite form result on uniform distribution is crucial here since the modulo $e_{2}(d)$ could be significantly greater than the range of $k_{1}$ (see Eq.
-(9)), which is different to the usual researches on Piatetski-Shapiro like sequences.
-Throughout the proof, $\mathcal{S}(x)$ denotes the number of elements of $\mathcal{S}$ not exceeding $x$.
+Throughout the proof, $\mathcal{S}(x)$ denotes the number of elements of $\mathcal{S}$ not exceeding $x$ and $x$ is supposed to be sufficiently large.
 Proof of Theorem 1.
@@
 For the innermost sum in the right–hand side above, we separate it into two cases according to $a_{1}=a_{2}$ or not. The contribution of the sum for $a_{1}=a_{2}$ can be controlled by
-$A(x)\mathcal{P}(x)\ll x$
-via Eq. (2) and the prime number theorem. Thus, to complete our proof, it would suffice to prove that
+$\mathcal{A}(x)\mathcal{P}(x)\ll x$
+via Eq. (2) and the prime number theorem. Thus, to complete our proof, it suffices to prove that
 $S(x):=\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}p_{2}<p_{1}\leq x\\ p_{1}-p_{2}=a_{2}-a_{1}\end{subarray}}1\ll x.$ (3)
@@
 $h_{i}=2^{\left\lfloor k_{i}^{r_{i}}\right\rfloor},~{}~{}~{}\widetilde{h_{i}}=2^{\left\lfloor k_{i}^{r_{i}}\right\rfloor}\quad(1\leq i\leq s-1),\quad\quad h_{s}=2^{\left\lfloor\lfloor k_{s}^{\lambda}\rfloor^{r_{s}}\right\rfloor},~{}~{}~{}\widetilde{h_{s}}=2^{\left\lfloor\left\lfloor\widetilde{k_{s}}^{\lambda}\right\rfloor^{r_{s}}\right\rfloor}.$
-For a given $g\leq x$, suppose first that there is at most one solution $k_{1}$ to the congruence equation
-$2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv g\ ({\rm{mod}}\ d),$ (7)
-then for those $d$ the double sums in right–hand side of Eq. (2) are bounded by
-$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\leq\prod_{2<p<\log x}\left(1+\frac{1}{p}\right)\ll\log\log x$ (8)
-which is admissible for our purpose, where the last estimate follows from the famous theorem of Mertens.
-We are left over to consider the remaining $d$ for which there are at least two solutions $k_{1}$ satisfying Eq.
-(7).
+For some integers $d$, suppose first that the number of $k_{1}$ satisfying the congruence
+$2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv g\ ({\rm{mod}}\ d)$ (7)
+is not large than $(\log x)^{1/r_{1}}/\log\log x$, then it is admissible for those $d$ since the double sums in right–hand side of Eq. (2) are bounded now by
+$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\frac{(\log x)^{1/r_{1}}}{\log\log x}$ $\leq\frac{(\log x)^{1/r_{1}}}{\log\log x}\prod_{2<p<\log x}\left(1+\frac{1}{p}\right)$
+$=\frac{(\log x)^{1/r_{1}}}{\log\log x}\exp\left(\sum_{2<p<\log x}\log\left(1+\frac{1}{p}\right)\right)$
+$<\frac{(\log x)^{1/r_{1}}}{\log\log x}\exp\left(\sum_{2<p<\log x}\frac{1}{p}\right)$
+$\ll(\log x)^{1/r_{1}},$ (8)
+where the last estimate follows from the famous theorem of Mertens.
+We are left over to consider the remaining $d$ for which the number of $k_{1}$ satisfying Eq.
+(7) is greater than $(\log x)^{1/r_{1}}/\log\log x$.
 In this case, let $k_{1,0}$ be such a solution.
@@
 $2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv 2^{\left\lfloor k_{1,0}^{r_{1}}\right\rfloor}\ ({\rm{mod}}\ d).$
-So we would get
+So, we would get
 $\left\lfloor k_{1}^{r_{1}}\right\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))$ (9)
-for some $10\leq\ell\leq e_{2}(d)-$ 1, where $e_{2}(d)$ is the least positive integer $m$ so that $d|2^{m}-1$. It follows that
-$e_{2}(d)f+\ell\leq k_{1}^{r_{1}}<e_{2}(d)f+\ell+1$
-for some $f\in\mathbb{Z}$ from Eq. (9) and vice versa, which clearly means that
-$\frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}<\frac{\ell+1}{e_{2}(d)},$
-where $\{y\}$ stands for the fractional part of $y$. It would be more pleasant to relax the above inequality to
-$\frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{\ell+1}{e_{2}(d)}$ (10)
-and we shall consider Eq.
-(10) from now on.
-We are in a position to employ a ‘finite’ form regarding the classical theorem of Weyl [24] on uniform distribution which was later established by Erdős and Turán [12, Theorem III].
-Lemma 1. If $\varphi_{1},\varphi_{2},\cdot\cdot\cdot,\varphi_{n}$ are real satisfying that
-$\bigg{|}\sum_{v=1}^{n}e^{ki\varphi_{v}}\bigg{|}\leq\psi(k)\quad(k=1,2,...,m,...),$
-then for $m\geq 1$ and arbitrary $0\leq\alpha<\beta\leq 2\pi$ we have
-[…]
-with a numerical constant $C$.
-Now we take $\varphi_{v}=2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\},n=\left\lfloor\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\right\rfloor,\alpha=\frac{2\pi\ell}{e_{2}(d)}$ and $\beta=\frac{2\pi(\ell+1)}{e_{2}(d)}$ in Lemma 1, then there is some absolute constant $C$ so that
-$\Bigg{|}\sum_{\begin{subarray}{c}k_{1}\leq\left(\log 2x/\log 2\right)^{1/r_{1}}\\ \frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{(\ell+1)}{e_{2}(d)}\end{subarray}}\!\!\!\!1-\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{e_{2}(d)}\Bigg{|}<C\left(\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{m+1}+\sum_{k=1}^{m}\frac{\psi(k)}{k}\right),$ (11)
-where $m\geq 1$ is a parameter to be decided later and $\psi(k)$ is the quantity satisfying
-$\bigg{|}\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\}}\bigg{|}\leq\psi(k).$ (12)
-To give an appropriate estimate of $\psi(k)$, we use the van der Corput type bounds [23, Theorems 5.9, 5.11 and 5.13].
+for some $10\leq\ell\leq e_{2}(d)-$ 1, where $e_{2}(d)$ is the least positive integer $m$ so that $d|2^{m}-1$.
+Furthermore, in case that (9) has more than $(\log x)^{1/r_{1}}/\log\log x$ solutions of $k_{1}$, we claim that
+$e_{2}(d)\ll_{r_{1}}(\log x)^{1-\frac{1}{r_{1}}}\log\log x,$ (10)
+which would be one of the key points in the following arguments.
+In fact, suppose that $L$ is the minimal distance between two consecutive solutions $k_{1}^{\prime}$ and $k_{1}^{\prime}+L$ of the congruence equation (9) with $k_{1}^{\prime}>(\log x)^{\frac{1}{2r_{1}}}$, then the number of solutions $k_{1}$ to (9) is not more than
+$(\log x)^{\frac{1}{2r_{1}}}+\frac{(\log x)^{1/r_{1}}}{L}+1.$
+Hence, by our assumption on the number of the solutions, we have
+$(\log x)^{\frac{1}{2r_{1}}}+\frac{(\log x)^{1/r_{1}}}{L}+1\geq\frac{(\log x)^{1/r_{1}}}{\log\log x},$
+which clearly means that $L\leq 2\log\log x.$ By the definition of $L$, we also have
+$\lfloor(k_{1}^{\prime}+L)^{r_{1}}\rfloor-\lfloor(k_{1}^{\prime})^{r_{1}}\rfloor\geq e_{2}(d),$
+from which it follows that
+$e_{2}(d)\leq(k_{1}^{\prime}+L)^{r_{1}}-(k_{1}^{\prime})^{r_{1}}+1\ll_{r_{1}}(k_{1}^{\prime})^{r_{1}-1}L\ll(\log x)^{1-1/r_{1}}\log\log x.$
+We now conclude from (2), (9), (10) and the above proofs that we only need to show
+$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1\ll(\log x)^{\frac{1}{r_{1}}}.$ (11)
+For any $k_{1}\leq(\log x)^{1/r_{1}}$ satisfies (9), we clearly have
+$e_{2}(d)q+\ell\leq k_{1}^{r_{1}}<e_{2}(d)q+\ell+1$
+for some nonnegative integer $q$, which is equivalent to
+$\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}<q\leq\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}.$ (12)
+It is plain that an integer $q$ satisfies (12) if and only if
+$\left\lfloor\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}\right\rfloor-\left\lfloor\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}\right\rfloor=1,$
+from which it follows that
+$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1$ $\leq\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\bigg{(}\left\lfloor\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}\right\rfloor-\left\lfloor\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}\right\rfloor\bigg{)}$
+$=\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\bigg{(}\frac{1}{e_{2}(d)}-\psi\left(\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}\right)-\psi\left(\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}\right)\bigg{)}$
+$\ll_{r_{1}}\frac{(\log x)^{1/r_{1}}}{e_{2}(d)}+O\bigg{(}\max_{|\theta|\leq 1}\bigg{|}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)\bigg{|}\bigg{)},$ (13)
+where $\psi(t)=\{t\}-1/2$ is the sawtooth function. Now, we write
+$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)=S_{1}+S_{2},$ (14)
+where
+$S_{1}=\sum_{k\leq e_{2}(d)^{1/r_{1}}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)\quad\text{and}\quad S_{2}=\sum_{e_{2}(d)^{1/r_{1}}<k\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right).$
+Trivially, we have
+$S_{1}\ll e_{2}(d)^{1/r_{1}}.$ (15)
+We are now in a position to use the following van der Corput type bounds [14, Lemma 4.3] for a suitable estimate of $S_{2}$.
 Lemma 1.
-Let $f(x)$ be real and have continuous derivates up to the $j$–th order, where $j\geq 2$.
-Let $\lambda_{j}\leq f^{(j)}(x)\leq h\lambda_{j}$ (or the same for $-f^{(j)}(x)$).
-Let $b-a\geq 1,J=2^{j-1}$.
-Then
-$\sum_{a<n\leq b}e^{2\pi if(n)}=O\left(h^{2/J}(b-a)\lambda_{j}^{1/(2J-2)}\right)+O\left((b-a)^{1-2/J}\lambda_{j}^{-1/(2J-2)}\right),$
-where the constants implied by big-$O$ are independent of $j$.
-Let $f(x)=\frac{kx^{r_{1}}}{e_{2}(d)}$ and $r$ be the least integer greater than $r_{1}$. Then
-$f^{(r)}(x)=\frac{k}{e_{2}(d)}r_{1}(r_{1}-1)\cdot\cdot\cdot(r_{1}-r+1)x^{r_{1}-r}.$
-Considering with $a\leq x<b\leq 2a$, we clearly have
-$\frac{kC_{1}}{e_{2}(d)(2a)^{r-r_{1}}}\leq f^{(r)}(x)\leq 2^{r-r_{1}}\frac{kC_{1}}{e_{2}(d)(2a)^{r-r_{1}}},$
-where $C_{1}=r_{1}(r_{1}-1)\cdot\cdot\cdot(r_{1}-r+1)$. Employing Lemma 2 we obtain that
-$\sum_{a<n\leq b}e^{2\pi i\frac{kn^{r_{1}}}{e_{2}(d)}}$ $\ll_{r_{1}}a\left(\frac{k}{e_{2}(d)a^{r-r_{1}}}\right)^{1/(2R-2)}+a^{1-2/R}\left(\frac{k}{e_{2}(d)a^{r-r_{1}}}\right)^{-1/(2R-2)}$
-$\ll a^{1-\frac{r-r_{1}}{2R-2}}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+a^{1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}},$
-where $R=2^{r-1}$. The dichotomy of $a$ between the interval $\left[1,\left(\log 2x/\log 2\right)^{1/r_{1}}\right]$ yields
-$\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\}}$ $=\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\frac{v^{r_{1}}}{e_{2}(d)}}$
-$\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}}.$
-Thus, by Eq. (12) we can choose $\psi(k)$ to satisfy that
-$\psi(k)\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}},$
-from which it follows directly via partial summations that
-$\sum_{k=1}^{m}\frac{\psi(k)}{k}\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{m^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}e_{2}(d)^{\frac{1}{2R-2}}.$ (13)
-We take $m=\lfloor\log\log x\rfloor$ and then insert Eq. (13) into Eq. (11). Then
-$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{(\ell+1)}{e_{2}(d)}\end{subarray}}\!\!\!\!1=\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{e_{2}(d)}+O_{r_{1}}\left(\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{\log\log x}+\sum_{k=1}^{\lfloor\log\log x\rfloor}\frac{\psi(k)}{k}\right).$ (14)
-Collecting Eqs. (2), (8), (13) and (14), we conclude from the analysis above that
-$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}d|a_{2}-a_{1}\\ 2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\!\!\frac{\mu^{2}(d)}{d}\ll_{r_{1}}(\log x)^{2-\frac{1}{r_{1}}}\log\log x+S_{1}(x)+S_{2}(x)+S_{3}(x)+S_{4}(x)$ (15)
+For any real number $t,$ let $\psi(t)=\{t\}-1/2$.
+Suppose $K\geq 3$ is a real number, $Y>0.$ Suppose $\alpha>0~{}(\alpha\notin\mathbb{Z})$ is a fixed real number.
+Then for any exponent pair $(\kappa,\lambda)$ we have
+$\sum_{K\leq k<2K}\psi\left(Yk^{\alpha}+\theta\right)\ll\frac{K^{1-\alpha}}{Y}+Y^{\frac{\kappa}{1+\kappa}}K^{\frac{\lambda+\kappa\alpha}{1+\kappa}}.$
+For $e_{2}(d)^{1/r_{1}}\leq K\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}$, by Lemma 1 with $Y=\frac{1}{e_{2}(d)}$ we have
+$\sum_{K\leq k_{1}<2K}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)$ $=\sum_{K\leq k<2K}\psi\left(Yk_{1}^{r_{1}}\right)$
+$\ll\frac{1}{YK^{r_{1}-1}}+Y^{\frac{\kappa}{1+\kappa}}K^{\frac{r_{1}\kappa+\lambda}{1+\kappa}}$
+$\ll\frac{e_{2}(d)}{K^{r_{1}-1}}+e_{2}(d)^{-\frac{\kappa}{1+\kappa}}K^{\frac{r_{1}\kappa+\lambda}{1+\kappa}}$
+since $Y=1/e_{2}(d)$, which combining a splitting argument yields
+$S_{2}=\sum_{e_{2}(d)^{1/r_{1}}<k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}}\psi\left(\frac{k^{r_{1}}}{e_{2}(d)}+\theta\right)\ll(\log x)^{\frac{1}{r_{1}}\cdot\frac{r_{1}\kappa+\lambda}{1+\kappa}}e_{2}(d)^{-\frac{\kappa}{1+\kappa}}+e_{2}(d)^{1/r_{1}}.$ (16)
+Taking (15) and (16) into (14), we have
+$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1\ll\frac{(\log x)^{1/r_{1}}}{e_{2}(d)}+(\log x)^{\frac{1}{r_{1}}\cdot\frac{r_{1}\kappa+\lambda}{1+\kappa}}e_{2}(d)^{-\frac{\kappa}{1+\kappa}}+e_{2}(d)^{1/r_{1}}.$
+Therefore, we have
+$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1\ll_{r_{1}}\mathcal{W}_{1}+\mathcal{W}_{2}+\mathcal{W}_{3},$ (17)
 where
-$S_{1}(x):=(\log x)^{2}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{de_{2}(d)},\qquad S_{2}(x):=\frac{(\log x)^{2}}{\log\log x}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d},$
-$S_{3}(x):=(\log x)^{2-\frac{r-r_{1}}{r_{1}(2R-2)}}(\log\log x)^{{}^{\frac{1}{2R-2}}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{de_{2}(d)^{\frac{1}{2R-2}}}$
+$\mathcal{W}_{1}=\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\frac{(\log x)^{1/r_{1}}}{e_{2}(d)},$
+$\mathcal{W}_{2}=\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}(\log x)^{\frac{1}{r_{1}}\cdot\frac{r_{1}\kappa+\lambda}{1+\kappa}}e_{2}(d)^{-\frac{\kappa}{1+\kappa}}$
 and
-$S_{4}(x):=(\log x)^{2-\left(\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}\right)}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)e_{2}(d)^{\frac{1}{2R-2}}}{d}.$
+$\mathcal{W}_{3}=\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}e_{2}(d)^{1/r_{1}}.$
 By a result of Erdős and Turán [12], we have
 $\sum_{2\nmid d}\frac{1}{d(e_{2}(d))^{\varepsilon}}<\infty$ (18)
-for any $\varepsilon>0$. Thus, we deduce that $S_{1}(x)\ll(\log x)^{2}$ and
-$S_{3}(x)\ll(\log x)^{2-\frac{r-r_{1}}{r_{1}(2R-2)}}(\log\log x)^{{}^{\frac{1}{2R-2}}}\ll(\log x)^{2}.$
-By Eq. (8), we also have $S_{2}(x)\ll(\log x)^{2}$. For $2\nmid d$, we have
-$2^{\varphi(d)}\equiv 1\ ({\rm{mod}}\ d)$
-by the Euler theorem, where $\varphi(\cdot)$ is the Euler totient function. So $e_{2}(d)\leq\varphi(d)\leq d$, from which it follows that
-$S_{4}(x)\leq(\log x)^{2-\left(\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}\right)}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d^{1-\frac{1}{2R-2}}}.$
-Changing back the sum to product, we get
-$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d^{1-\frac{1}{2R-2}}}\leq\prod_{2<p<\log x}\left(1+\frac{1}{p^{1-\frac{1}{2R-2}}}\right)=\exp\left(\sum_{2<p<\log x}\log\left(1+\frac{1}{p^{1-\frac{1}{2R-2}}}\right)\right).$
-Employing the elementary inequality $\log(1+z)<z$ for any $z>0$, the right–hand side above is
-$<\exp\left(\sum_{2<p<\log x}\frac{1}{p^{1-\frac{1}{2R-2}}}\right).$
-Using the prime number theorem as well as the partial summations, we have
-$\sum_{2<p<\log x}\frac{1}{p^{1-\frac{1}{2R-2}}}\ll\frac{(\log x)^{\frac{1}{2R-2}}}{\log\log x},$
-which implies $S_{4}(x)\ll(\log x)^{2}$ on noting that
-$\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}>\frac{1}{2R-2}.$
-Therefore, the right–hand side of Eq. (15) is bounded by $(\log x)^{2}$ which proves Eq. (5) and hence Eq. (3). ∎
+for any $\varepsilon>0$. Using (18), we know that
+$\mathcal{W}_{1}\leq(\log x)^{1/r_{1}}\sum_{\begin{subarray}{c}2\nmid d\end{subarray}}\frac{1}{de_{2}(d)}\ll(\log x)^{1/r_{1}}$
+and together with the estimate given in (2), we get
+$\mathcal{W}_{3}$ $\ll\Big{(}(\log x)^{1-1/r_{1}}\log\log x\Big{)}^{1/r_{1}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}$
+$\ll(\log x)^{\frac{1}{r_{1}}-\frac{1}{r_{1}^{2}}}(\log\log x)^{\frac{1}{r_{1}}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}$
+$\ll(\log x)^{\frac{1}{r_{1}}-\frac{1}{r_{1}^{2}}}(\log\log x)^{\frac{1}{r_{1}}+1}$
+$\ll(\log x)^{\frac{1}{r_{1}}}.$
+On noting (17), to establish the estimates of (11) it remains to show $\mathcal{W}_{2}\ll(\log x)^{\frac{1}{r_{1}}}$.
+Again, by (18) and the estimates in (2), it suffices to choose a suitable exponent pair $(\kappa,\lambda)$ such that
+$\frac{r_{1}\kappa+\lambda}{1+\kappa}<1.$ (19)
+For any integer $q\geq 1$, we know that
+$\left(\frac{1}{4Q-2},1-\frac{q+1}{4Q-2}\right)$
+is an exponent pair, thanks to the display of [14, Page 60], where $Q=2^{q}$. We now take $q=\lfloor r_{1}\rfloor+1$ and then
+$r_{1}\kappa+\lambda=\frac{r_{1}}{4Q-2}+1-\frac{q+1}{4Q-2}<1,$
+[…]
 Acknowledgments
-The author is supported by National Natural Science Foundation of China (Grant No.
+We thank Rui-Jing Wang for his interests and pointing out some inaccuracies in the arguments of the former version.
+Yuchen Ding is supported by National Natural Science Foundation of China (Grant No.
 12201544), Natural Science Foundation of Jiangsu Province, China (Grant No.
@@
 2022M710121).
+Wenguang Zhai is supported by National Natural Science Foundation of China (Grant Nos. 12471009, 12301006) and by Beijing Natural Science Foundation (Grant No. 1242003).